### v1

# code-for-first-row = , code-for-last-row = , code-for-first-col = , code-for-last-col =

Abstract
Posets are discrete mathematical structures which are ubiquitous in a broad range of data analysis and machine learning applications. Research connecting posets to the data science domain has been ongoing for many years. In this paper, a comprehensive review of a wide range of studies on data analysis amd machine learning using posets are examined in terms of their theory, algorithms and applications. In addition, the applied lattice theory domain of formal concept analysis will also be highlighted in terms of its machine learning applications.
Keywords— posets, machine learning, data analysis, formal concept analysis, lattice theory
1 Introduction
Data analysis plays a pivotal role in today’s data-driven world. Machine learning is used to automate the data analysis process, and enables the the extraction of valuable insights by making accurate predictions from large complex datasets. Data often have complex structure which can usually be endowed with a natural order. Such complex data can be derived from text, chemical structures, biological structures, images, videos and so on. Mathematical techniques and concepts serve as powerful tools in the development of new algorithms and methods for machine learning and data analysis. For some time now, many studies based on the use of partially ordered sets(posets) for machine learning and data analysis have emerged. Furthermore, over this time the literature on these topics has significantly expanded in volume, scope, and the range of its applications in various fields. However, no recent surveys exist to collect and organize this knowledge thereby impeding the ability of researchers and machine learning scientists alike to utilize it for various data science or artificial intelligence tasks. In order to address this gap, we present an up-to-date overview with an emphasis on the most prominent and currently relevant works on data analysis and machine learning methods in which the role of posets has been established. By covering earlier works as well as more recent advances, our survey aims to provide researchers, applied mathematicians, data scientists, statisticians, social scientists and practitioners new to the field as well as all those aiming at keeping pace with innovation in this interesting, growing and promising research field with a solid understanding of the different conceptual and methodological approaches which incorporates posets into data analytics or machine learning applications.
The rest of this survey is organized as follows. Section 2 introduces and outlines the key concepts on partially ordered sets and the related field of lattice theory. Section 3 presents a summary of various research studies on machine learning and deep learning with respect to poset theory, lattice theory and formal concept analysis. Section 4 focuses on cluster analysis using posets and formal concept analysis methods. Sections 5-7 presents a summary on data analysis from a poset theoretical perspective covering its multidimensional, exploratory and descriptive aspects respectively. Section 8 presents a collection of applications across various domains including software packages, datasets and algorithms for this purpose. Section 9 discusses the current challenges and suggests future directions. Section 10 summarizes the paper.
2 Basics of Partial Order Theory
[…]
  1. 1.
    reflexive: $x\in X,x\preccurlyeq x$,
  2. 2.
    antisymmetric: $x,y\in X$ if $x\preccurlyeq y$ and $y\preccurlyeq x\implies x=y,$
  3. 3.
    transitive: $x,y,z\in X$ if $x\preccurlyeq y$ and $y\preccurlyeq z\implies x\preccurlyeq z.$
[…]
  • Maximal elements or objects of the poset $P$ are the set of elements or objects of $X$ for which no other object or element $y$ satisfies the relation $y>x.$ If $x$ is the only maximal object or element then it is referred to as the ”greatest” object or element. In a totally ordered set, the terms maximal element or object and greatest element or object coincide.
  • Minimal elements or objects of the poset $P$ are the set of elements or objects of $X$, for which no other object or element $y$ exist such that $y<x$. If $x$ is the only minimal object or element then $x$ is referred to as the “least” object or element. In a totally ordered set, the terms minimal element or object and least element or object coincide.
  • Chain is a subset C of X, where any object or eleement is mutually comparable with all other objects or elements of C. That is, a chain is a poset such that no element or object can be added between any two of its elements without losing the property of being totally ordered.
  • Antichain is a subset $C^{\prime}$ of X, where each object or element of $C^{\prime}$ is mutually incomparable with all other objects or elements in $C^{\prime}$. That is, all the objects or elements of the set $C^{\prime}$ are pairwise incomparable elements and as such they can never have the property of being totally ordered.
  • A cover relation for ${\mathcal{P}}=(X,\preccurlyeq)$ is the set of pairs $(x,y)$ such that $y\in X$ covers $x$ whenever there exist no element $z\in X$ such that $x<z<y.$
Partially ordered sets can often be graphically visualized using Hasse diagrams[62]. A Hasse diagram can be derived from a directed acyclic graph, where the vertices are representing the objects and a line relates object $x$ with $y$ whenever $x<y$. In the case of a transitivity relation whenever $x<y$ and $y<z$ then it suffices to draw a line only for $x<y$ and $y<z$ and not for $x<z$. Hasse diagrams can be viewed as a powerful tool for unifying ideas and concepts. Hasse diagrams reveal an intricate network of comparabilities and incomparabilities, maximal and minimal elements. In addition, Hasse diagrams are also characterized with a structure that comprises levels, chains and antichains.
Hasse diagram: 2431
Linear Extensions: 231432143412
Finite posets can also be represented by square matrices which comprise: incidence [262, 262], cover [263], and mutual ranking probability matrices[264]. The mutual ranking probability matrices are a class of matrices which comprise and convey information on the dominance among statistical units All the three main types of poset matrices are useful in conveying essential information on the structure of the order relation from different perspectives.
Posets can also be subdivided into variour classes such as series-parallel posets, semantic posets and ranking posets etc. Series-parallel posets are characterized by the N-free partial orders. Series-parallel partial orders have been applied machine learning of event sequencing in time series data [134]. Semantic posets which are used for compositional generalization in semanti parsing [105]. The ranking poset $\mathfrak{M}_{n}$[129] is the poset in which the elements are all possible cosets $\mathfrak{G}_{\lambda}\pi$, where $\lambda$ is an ordered partition of $n$ and $\pi\in\mathfrak{G}_{n}.$ The ranking poset may be useful in model selection and validation in machine learning.
Posets are natural models for many statistical applications [219, 165]. Partial orders are the natural mathematical structure for comparing multivariate data that lack a natural order [143]. For instance, multivariate data on colours has no natural ordering and so the rank features in colour spaces are based on partial orders and this approach is generalizable to similar contexts [276]. It also plays a key role in decision making in environmental science and chemistry [33, 107] etc.
The study of partially ordered sets and lattices has been covered by an extensive amount of mathematical literature. Birkhoffs (1967) [258] and Grdtzer’s (1978)[259] books on lattice theory are considered to be classics. Lattices[181] are a broad structured class of posets. Lattice theory has shown its ability to contribute to solving problems in a wide variety of applications. Formally, a lattice [104] is a poset, in which every pair of elements has both a least upper bound and a greatest lower bound. In other words, it is a structure with two binary operations: join and meet. The following key sub-definitions summarize a lattice structure. A meet semilattice is a poset for which any two elements a and b have a greatest lower bound denoted $a\wedge b$. The greatest lower bound of a and b is the largest element that is still less than both of them. In a lattice, the greatest lower bound must be unique. The greatest lower bound of a and b is also called the meet or infimum of $a$ and $b$. A join semilattice is a poset for which any two elements $a$ and $b$ have a least upper bound, denoted $a\vee b$. The least upper bound of $a$ and $b$ is the smallest element that is still greater than both. In a lattice, the least upper bound must be unique. The least upper bound of $a$ and $b$ is also called the join or supremum of $a$ and $b$. If a poset is both a meet semilattice and a join semilattice, then the poset is also a lattice.
Example of a lattice : 2134
In the example above: The maximal elemnt is $4$ and the greatest element is $4$ Similarly, the minimal element is $1$ and the least element is $1$
3 Machine learning using posets and formal concept analysis
3.1 Overview
Machine learning generally involves the creation of knowledge from a training dataset by using models and algorithms to learn from data in order to make predictions, find pattrens, classify data or automate decision making processes. Machine learning has progressed significantly over the past two decades, and this has been driven by the development of new learning algorithms, theory, and high performance computing architecture [290]. There are many different kinds of machine learning algorithms. The most well-known ones are supervised, unsupervised, semi-supervised, and reinforcement learning. The goal of machine learning algorithm is to optimize the performance of a system when handling new instances of data through user defined programming logic for a given environment [301]. In order to build more efficient machine learning algorithms or models so as to improve on their performance and accuracy in most predictive task, various ensemble learning methods have been developed. Ensemble learning[291] is an approach which aggregates the predictions of two or more models fitted to the same data in order to minimize error. In the area of supervised learning, the most popular ensemble machine learning techniques are bagging, boosting and stacking, which use multiple learning algorithms or models to produce one optimal predictive model. Ensemble methods can broadly be categorized into sequential ensemble techniques and parallel ensemble techniques. Bagging aims to reduce variance by adopting parallel ensemble learning, boosting aims to reduce bias by adopting sequencial ensemble learning , and stacking aims to improve prediction accuracy by adopting parallel ensemble learning. Sequential ensemble techniques generate base learners in a sequence that are characterized by the dependence between the base learners. On the other hand the parallel ensemble technique base learners are generated in parallel in such a way that independence is enforced between the base learners. Dagging(Disjoint samples aggregating) is another useful type of parallel ensemble technique that was introduced in [300]. Dagging is similar to bagging, but instead of bootstrapping, it uses stratified sampling by creating a number of disjoint groups and stratified data from the original learning data set, with each considered as a subset of learning. In unsupervised learning, ensemble clustering aims at combining the outputs of several clustering algorithms to form a single clustering structure (crisp or fuzzy partition, hierarchy) [302, 303].
In recent years, deep learning[292] has been the most popular computational approach in the field of machine learning and has unique advantages when dealing with high-dimensional nonlinear problems. The performance of the machine learning algorithm relies heavily on the representation and finding a good representation can facilitate discovering structure in the input data by the learning algorithm. Deep learning as a model is a specific type of representation learning[283] that uses layered algorithms known as artificial neural networks, which attempts to mimic the human brain through a combination of data inputs, weights, and bias, to learn representations of data. It has succesfully been used in a wide range of disciplines such as disciplines, including cybersecurity, natural language processing, visual recognition, machine translation, robotics, recommendation systems etc. There are several types of deep learning algorithms. These are : convolutional Neural Network, Long Short Term Memory, GraphNeural Networks etc. Traditionally, machine-learning and deep learning algorithms have been used on data represented in euclidean space such as image data which can be represented as a regular grid of pixel values. On the other hand, graph data cannot be represented on euclidean space [307]. Graph Neural Networks (GNNs) are a class of deep learning methods designed to perform inference on data described by graphs which are non-euclidean (non-metric) structured. Posets are special classes of directed graphs. There is also a growing use of graph neural networks for the modeling and analysis of partially ordered data [99, 83, 105, 304, 306]. More recently, the problem of machine learning on meet/join lattices and posets was extensively studied in [180]. The main focus of the study was on developing methods that are based on generalized convolutions and sparse Fourier transforms algorithms (SFTs) that are capable of learning set functions. The first lattice-based machine learning models were introduced by V.K. Finn[295, 293, 294, 296]. It was based on a closure systems JSM-method1 of automated hypothesis generation. In this model positive hypotheses are searched among intersections(similarity as meet operations) of positive example descriptions (object intents), likewise for negative hypotheses [297]. The JSM method also identifies data patterns by means of induction and it can also be used as a logical rule-based classification method formulated in terms of formal concepts [298]. Machine learning has been studied from many perspectives in the context of poset theory and lattice theory. These include : Machine Learning Performnce Comparison[25, 129], Multilabel Classification [76], Ensemble Classification[129], Sequential Clasification [134, 166, 166, 165, 172]; Natural Language Processing [83, 132, 105], Deep Unsupervised Learning [20], Semi-Supervised Learning[99], Learning to Rank[270], Time Series Modelling [288, 134] .
3.2 Comparing the performance of machine learning algorithms using posets
Performance comparison of different learning algorithms is a very common task in the field of machine learning. The evaluation of several machine learning algorithms based on the concept of depth functions was recently introduced in [25]. Depth functions [218] were originally formulated to extend the univariate notions of median, quantiles, ranks, signs, and order statistics to the setting of multivariate data. Blocher & Schollmeyer [25] uses an adaptation of the well-known simplicial depth and then applied it to the set of all partial orders referred to as the union-free generic (ufg) depth. The purpose is to develop a framework to show how samples of poset-valued random variables can be descriptively analyzed by utilizing a generalized concept of data depth based on the concept of depth functions. A key property of the ufg depth is that it considers partial orders as a whole, and not limited to pairwise comparisons since as the procedure for computing the depth of a partial order cannot be broken down using simple sum statistics, but requires the partial order as a holistic entity. The ufg depth for a comparison of machine learning algorithms is based on multidimensional performance measures. It works by comparing the various algorithms not with respect to one unidimensional criterion such as balanced accuracy, but instead it considers a whole set of performance measures.
3.3 Natural Language Processing
Compositionality[220] which describes the ability to easily recombine familiar units like words into novel phrases and sentences has been the focus of intense interest in artificial intelligence in recent years. However, existing neural network based models have proven to be extremely deficient in such a capability[308]. In order to improve on the compositional generalization ability of neural encoder-decoder architectures for natural language understanding, Guo et al. [105] introduce a hierarchical poset decoding paradigm with the goal of taking into consideration the partial permutation invariance and the poset structure of semantics. Compositional generalization is important in tasks that can involve complex natural language utterances, which can be associated to many equivalent meaning representations. One of such task is dependency parsing. Semantic dependency parsing[221] is the task of mapping natural language sentences into representations of their meaning in the form of directed graphs on words. Dependency parsing is generally used to identify semantic relations between words in a sentence by analyzing their grammatical structure. Dyer [83] proposes a novel method of for converting dependency trees to surface orders via syntactic word embeddings and edge-weighted posets. Graph neural network were employed to learn the edge weights of the poset representing the surface order and the training was carried out on the Universal Dependencies (UD) corpora. [222]
3.4 Classification
Most major studies in artificial intelligence has concentrated significantly on developing mechanisms for classification. Most prediction tasks in machine learning are considered as classification when the target variable is discrete. It is common to find learning tasks that involve ranking of objects particularly in domains that generate relational data where the goal of inferring is conditioned on a particular target object . Classification is the task of associating a single label $y\in Y$ with a covariate $x$. A generalization of this problem is conditional ranking, the task of assigning to $x$ a full or partial ranking of the items in Y. Lebabnon and Lafferty[129] propose a unifying algebraic framework for ensemble methods of classification and conditional ranking, focusing on the case of discrete and ranked inputs. Their approach is based on the ranking poset having a fixed size of items, which consists of the collection of all full and partial rankings equipped with the partial order given by refinement of rankings. The structure of the poset of partial ranking over Y gives rise to natural invariant distance functions that generalize Kendall’s Tau[309] and the Hamming distance. Using the invariance properties of the distances, a generative interpretation of the probabilistic model is derived, and this is suggested to be useful in model selection and validation[129]. Zhang [152] presents a framework for ranking with classification rules. Classification rules can be generated using inductive learning algorithms on data [226]. Sahami [152] introduces an inductive machine learning algorithm called Ruleamer which roduces a set of symbolic classification rules as output using lattices.The Ruleamer algorithm takes as input (i) a lattice, L, (ii) a set of instance classification labelings, C, which correspond to the instance nodes in L, and (iii) a noise parameter, N, indicating a percentage by which each induced rule can misclassify some portion of the training instances to which it applies. Learning classification rules has also been extended to sequence data [224, 225]. The problem of sequential classification is considered when classification states form lattices and response distributions are experiment-specific. [172] proposes a Bayesian framework for sequential classification on finite lattice models in which response distributions are allowed to vary according to experiment When the models are latent and complex, such as in cognitive applications, it becomes imperative to have available a variety of data analytic tools for fitting the models, and for the validation of assumptions that are made regarding the specification of class conditional response distributions for experiments. A framework for model fitting and analysis of experiments has been developed byTatsuoka[165]. This led to the development of a general theorem in [166, 167] on the asymptotically optimal sequential selection of experiments which is presented and applied to a Bayesian classification problem when the parameter space is a finite partially ordered. In general, lattices can be used as classification models when it is natural to consider classification states as partially ordered. Examples include cognitive modeling, such as in neuropsychological (NP) assessment, educational testing, and the group testing data.
3.5 Deep Unsupervised Learning
Deep unsupervised learning[310] is an emerging discipline in the field of artificial intelligence and machine learning. It can serve as an alternative to supervised approaches in representation learning. In particular, unlabelled data is much cheaper to obtain which is a key advantage when adopting deep unsupervised learning for complex processing tasks involving large datasets. One of such task is visual similarities learning[227] in the field of computer vision. Various approaches have been adopted for visual similarity learning such as the use of convolution neural networks(CNN) [228]. But this approach has a major limitation in that it requires millions of samples of supervised training in order to achieve a good performance. Bautista et al.[20], propose an unsupervised approach to similarity learning based on CNNs by framing it as a combination of surrogatee (i.e. artificially created) classification tasks and poset ordering. This approach which combines classification with a partial ordering of samples solved some of the major problems of visual similarity learning such as:
  • a large number of training samples are not assigned to any of the compact surrogate classes since their mutual similarities or dissimilarities cannot be easily established;
  • inability to perform joint optimization of the different classification tasks resulting to mutually conflicting relationships since transitivity cannot be established ;
  • expensive cost or difficulty in obtaining a large number of samples for training.
3.6 Semi-Supervised Learning
Semi-supervised learning[229] is a branch of machine learning that bridges supervised and unsupervised learning by combining both labeled and unlabeled data to train artificial intelligence (AI) models for classification or regression. Semi-supervised classification methods are particularly relevant to scenarios where labelled data is scarce, expensive or difficult to obtain. Graph-based semi-supervised learning methods[230] have demonstrated their advantages in various domains due to their uniqueness of structure, the universality of applications, and their scalability to large scale data. Graph-based learning methods explicitly consider the relations between two entities (i.e.vertices). In an ordinary graph an edge connects exactly two vertices. On the other hand, a hypergraph represents a generalization of a graph in which an edge can join any number of vertices. But in most settings hypergraph structure represent the hyperedge as an unordered set of vertices without taking into consideration the possibility of ordering relationships among vertices which is inherent in most real-world data. This motivated Feng et al. [99] to suggest two related solutions which are:
  • the construction of a new data structure named Partial-Order Hypergraph, which specifically injects the partially ordering relations among vertices into a hyperedge ;
  • the development of regularization based learning theories for partial-order hypergraphs which generalizes conventional hypergraph learning by incorporating logical rules that encode the partial-order relations.
In particular, graph gonvolution network(GCN)[312] was applied in a semi-supervied setting for learning over the partial-order hypergraphs in [99].
3.7 Time Series Modeling
Many machine learning and statistical models have been suggested for time series modeling of sequential data [313, 314]. One way of generating sequence data is from a series of discrete events that occur over time. Event sequence data arise in many applications, such as web browsing, e-commerce, and monitoring of processes. An important problem in mining sets of sequences of events is to get an overview of the ordering relationships in the data. [134] introduce a method for discovering partial orders from a sequence of events. The main idea of the method is to view a partial order as a generative model for sequences and by describing a set of sequences by means of mixture models of partial orders. In addition, it is also shown that the likelihood of a given partial order is inversely proportional to the number of total orders compatible with the partial order. The computation of the number of linear extensions was restricted to series paprallel posets which are computationally efficient.
Nicholls et al. [288] studies Bayesian inference for partial orders from random linear extensions using time-series data of social hierarchy during the time period duringthe 12th century. The outcome of the approach a new class of models for rank-order time series data in which actors are listed in order of precedence are presented as a collection of lists. The lists are modeled as a realisation of a queue in which queue-position is constrained by an underlying social hierarchy corresponding to a partial order in which the lists are random linear extensions. A consistent prior for the stochastic process of partial orders based on a latent variable representation for the partial order Model fitting to witness lists from Royal Acta from England, Wales and Normandy in the eleventh and twelfth centuries.
3.8 Learning to rank
Learning to rank (LTR)[315] describes a class of algorithmic techniques that apply supervised machine learning to solve ranking problems from a wide range of applications in domains such as information retrieval, recommendation systems, search engine optimization etc. It arises from the need to obtain information in the order of relevance during retrieval of a prediction task. Several methods have been proposed for LTR and they can be categorized or grouped into three main approaches: pointwise, pairwise , listwise. Pointwise approaches look at a single document at a time in the loss function. They essentially take a single document and train a classifier / regressor on it to predict how relevant it is for the current query. Pairwise approaches look at a pair of documents at a time in the loss function. Given a pair of documents, they try and come up with the optimal ordering for that pair and compare it to the ground truth. The goal for the ranker is to minimize the number of inversions in ranking i.e. cases where the pair of results are in the wrong order relative to the ground truth. Listwise approaches directly look at the entire list of documents and try to come up with the optimal ordering for it.
In [268], it is noted that not all rankings can follow a strict total order in the context of learning to rank. As such a proposed method of relaxing the conventional setting such that predictions are given in terms of partial instead of total orders is outlined. The key idea is that if a model is uncertain about the relative order of two alternatives, in which case it is unable to clearly determine whether the former should precede the latter and vice-versar, it may decline from making a decison and instead declare such pair of alternatives as incomparable. Such learning to ranking problem can be formulated as follows: Given a set of training instances $\{x_{1},...,x_{n}\}\subseteq\mathcal{X}$ and a set of labels $\mathcal{Y}=\{y_{1},...,y_{k}\}$ endowed with an order $y_{1}<y_{2}<,...,<y_{k}$ such that for each training instance $x_{l}$ it can be associated a label $y_{l}.$ The problem is to determine a ranking function that orders a new set of instances $\{x^{\prime}_{j}\}^{t}_{j=1}$ according to their (unknown) preference degrees. The performance meaures for this are: AUC($k=2$, for the bipartite ranking) and C-index in the polytomous case ($k>2$, for $k$-partite ranking). As now the ranker has the ability to reject predictions, there is a trade-off between correctness and completeness. Correctness is measured by gamma rank correlation[269] is a measure of rank correlation, i.e., the similarity of the orderings of the data when ranked by each of the quantities. Completeness measure punishes the abstention from comparisons that should actually be made. Furthermore, a preference relation[268] $P:A\times A\rightarrow\left[0,1\right]$ provides a measure of support for the pairwise preference $a\succ b$ with $P(a,b)=1-P(b,a)$ for all $a,b\in A$ which can be considered an application of a generic approach to transform every ranker into a partial ranker via ensembling such that
  1. 1.
    For any ranker $L$, train $k$ ranking models $M_{1}…M_{k}$ by resampling from the original data set, i.e., by $k$ bootstrap samples. By querying these models, $k$ rankings $\succ_{1}...\succ_{k}$ are generated.
  2. 2.
    For each pair of alternatives $a$ and $b$, the degree of preference is defined as: $P(a,b)=\frac{1}{k}\lvert\{i\mid a\succ_{i}b\}\rvert.$
3.9 Machine learning using formal concept analysis
Formal concept analysis(FCA) [7] is a well known method which stems directly from partial order and lattice theory for conceptual knowledge representation and data analysis in information science. Formal concept analysis was introduced in the seminal work of Wille[181] in 1982 as a result of an attempt to restructure mathematical order and lattice theory. The fundamental theory underlying FCA is the representability of complete lattices by ordered sets of their meet- and join-irreducibles [231]. Since ordered sets of irreducibles are naturally represented by binary matrices, this makes it possible to apply certain aspects of the lattice theory to the analysis of data given by object-attribute matrices.[198]. Furthermore, the row of the binary matrix is usually interpreted as an object and a column is interpreted as a binary attribute. If an object possesses an attribute the corresponding entry is 1 and if it does not possess it then the correesponding entry is 0 [298]. It has also been shown that through a process of discretising and booleanising data, it is possible to convert a variety of data sets into formal contexts or concept lattices [278, 279, 280]. FCA can be viewed as an unsupervised machine learning technique. In this direction, it takes an input binary relation which is represented by a binary matrix. Its goal is usually to determine the natural concepts described in the data, and then organizes the concepts in the form of a Hasse diagram corresponding to a partial ordering [70, 316, 317]. Due to its generality, simplicity, and powerful mathematical foundation, there is currently a growing adoption of FCA in a wide variety of data science tasks. Its advantage can be seen in the several FCA based machine learning methods that gave competitive results in various machine learning problems compared to classical methods [185]. In a recent study on the application of formal concept analysis in image classification problems, Khatri[184] establishes several advantages of formal concept analysis over convolutional neural networks for image classification task. These advantages are :
  • Firstly, FCA provides clarity which can be useful for interpretability of machine learning classification tasks. Furthermore, FCA facilitates examination of classification hierarchy by providing a visual representation of data in the form of a lattice.
  • Secondly, during the classification process, data can be added and removed from the lattice without the need for retraining. Moreso, the FCA classifier outperforms the Random forest classifier, which is a human interpretable model.
An important feature of FCA-based classification methods is that they do not make any assumptions regarding statistical models of a dataset[216]. A typical refinement of the idea of learning is learning from positive and negative examples, when a learning system obtains descriptions of positive and negative examples and constructs a generalization of positive examples that does not “cover” negative examples[198]. There are several FCA-based models for classification based on learning from positive and negative examples [187], [197],[198]. In [233] a genetic algorithm based machine learning method is presented for automatically learning object-oriented hierarchies very similar to lattice based concept hierarchies from large data set. Ikeda and Yamamoto[239] propose a method for classifying data that generates a concept lattice and selects appropriate formal concepts in the lattice. The method is useful for robust feature selection during classification by securing sufficient storage space in order to maintain both the selected concepts and redundant concepts. This is beneficial in terms of reduced time duration in solving practical problems related to a variety of large data sets. Formal concept analysis has been applied in the context of ensemble learning: Dagging [267, 256], Recommendatory-based Multiple Classifier System [266]
FCA based approaches have also been used for the automated discovery of domain-specific ontologies from textual descriptions of domain entities [318, 319]. Ontology learning[320] refers to the process of automatically extracting and constructing knowledge structures or models from unstructured or semi-structured data sources such as text, speech, images, or sensor measurements. These knowledge structures typically take the form of annotated taxonomies, concept hierarchies, or domain-specific ontologies that capture various aspects of the underlying domain or subject matter. Ontologies are a key element of the semantic web. They aim to capture basic knowledge by providing appropriate terms and formal relationships between them, so that they can be used in a machine-processable manner. Relational Concept Analysis (RCA) is an extension of the FCA framework to take into account multi-relational datasets, RCA generates a family of concept lattices, precisely one for each category of objects Using formal concept analysis as well as relational concept analysis different approaches have been implemented for ontology learning and ontology ex-traction [234, 235, 236, 237, 232]. Many machine learning based AI systems have been designed as black boxes. which makes it challenging in achieving interpretability in AI models due to the complexity of AI algorithms. Sangroya et al.[281] propose a general concept lattice theory based framework for explainability of AI models. The main idea is that given an outcome of a deep learning model and a domain ontology, the objective of their proposed solution is to identify an explanation that can point the user to the prominent feature set for a certain outcome.
4 Clustering Partially Ordered Data
Cluster analysis is a multivariate data analysis technique for identifying structural patterns and relationships within complex datasets with the goal of grouping similar data points into clusters. In the process, it attempts to discover the internal structure of a data set by considering the similarity between objects based on some type of attributes that the objects may possess in varying degrees. The ordinal model for clustering with posets was introduced in [116], where it is shown that the characterization of flat cluster methods[115] leads to a universal mapping problem in the theory of partially ordered sets. A persistent underlying theme involves generalized notions of adjoints of order preserving mappings between posets [117]. In addition, [117] attempts to develop a clustering scheme based on dissimilarities measured in posets. Many of the cluster methods that are used in the construction of classificatory systems operate on data in the form of a dissimilarity coefficient on a set of objects [21]. There are two methods in cluster analysis, namely hierarchical and non-hierarchical. Mathematically, hierarchical clustering involves the creation of a nested sequence of partitions of a set, whereas non hierarchical clustering deals with a single partition.The hierarchical clustering method consists of the agglomerative and divisive methods. Divisive clustering use a top-down approach to combine all the data points as a single cluster and then divide them as the distance between them increases. On the other hand, agglomerative clustering is a bottom-up approach which divides the data points into different clusters and then aggregates them as the distance decreases. Agglomerative clustering can be categorized into single linkage and complet linkage. Single linkage is a method that focuses on minimum distances or nearest neighbor between clusters while complete linkage focuses on maximum distance or furthest neighbor between clusters [285]. Sabara et al. [190] use the single linkage and complete linkage agglomerative hierarchical clustering method on the Hasse graph of posets in a multidimensional data analytics setting. Similarly, [189, 122] also develop methods that combine poset methodology with hierarchical clustering from a multidimensional data perspective. In particular, [189] develops a method of hierarchical stratification which is interpretable using the partial order Hasse graph and the evaluation process is such that it takes into consideration the characteristics of multi-dimensional indicators, division into several layers whereby different layers have different properties and characteristics.
Janowitz [118] presents a comprehensive review of clustering including through the use of lattices that generalize trees. Ontologies represent data relationships as hierarchies of possibly overlapping classes. Ontologies are closely related to clustering hierarchies.Dissimilarity is a common intermediary used by clustering methods to classify data. In [131], it is shown that modeling ontologies as posets over the subset relation enables classical clustering algorithms which take dissimilarity matrices as inputs always incorporate all available information, as such there is no loss of information. This served as the basis for the development of a clustering algorithm, that generates a partially ordered set of clusters from a dissimilarity matrix. Dissimilarity matrices simplify some of the problems associated with clustering high-dimensional datasets, since their size is only a function of the number of objects $\mathcal{O}(N^{2})$ , and independent of the objects’ dimensions.
Formal Concept Analysis (FCA) is an approach for identifying structure present in domains. The FCA can be seen as a conceptual clustering method. Michalski(1980)[321] introduced conceptual clustering as a new branch of machine learning. Since its introduction there has been significant interest in the different machine learning tasks that can utilize conceptual clustering methods. Work in Artificial Intelligence has concentrated significantly on developing mechanisms for classification, and the conceptual representations has been pivotal in supporting this effort. Conceptual clustering [69] is concerned with the problem of grouping unlabelled objects into classes. Three key features of conceptual clustering methods are:
  1. 1.
    Every output class, in addition to being characterized by its extensional description (i.e., the set of objects covered by the class), is also characterized by an intensional (conceptual);
  2. 2.
    Output classes (often referred to as concepts) are arranged into a hierarchy based on their generality/specificity
  3. 3.
    the process of class formation is incremental; that is, the processing of the $n$th object does not require extensive reprocessing of the $(n-1)$ objects already seen. description
The theory of concept or galois lattices provides a simple and formal approach to conceptual clustering. Carpineto[68] develop the clustering algorithm GALOIS which is able to determine the concept lattice corresponding to a given set of objects. It was demonstrated to be useful for class discovery and class prediction with a the time complexity of each update ranging from $O(n)$ to $O(n^{2})$ where n is the number of concepts in the lattice. Fisher [322]presented conceptual clustering as an extension of numerical taxonomy. In [240], the theoretical background of the application of Hierarchical Cluster Analysis HCA to improve on the understanding of Hasse Diagrams derived from the ranking process is presented. The use of HCA is based on the idea of reducing the number of elements in a poset by the selection of representatives from the original set used for ranking. As such the clusters derived from HCA can be interpreted as similarity classes in which one of its members (the nearest to the centre of the cluster) can be selected as representative of its class. Generally, incremental hierarchical conceptual clustering algorithms work by defining a set of hierarchy-change operators and carrying out a hill-climbing search over the space of possible concept hierarchies with the goal of determining a “good” hierarchy [214]. These algorithms have been implemented by several incremental hierarchical conceptual clustering systems have been developed over the years([199, 200, 201, 202, 203, 204, 210]). Although they differ in many respects (e.g., concept representation, hierarchy evaluation criteria, disjoint versus nondisjoint classes), they usually solve the problem by defining a set of hierarchy-change operators and carrying out a hill-climbing search through the space of possible concept hierarchies aimed at finding a “good” hierarchy (Gennari et al., 1989). These systems perform well in many experimental settings, but the criteria they use to modify a [69]. Another major developemnt in hierarchical conceptual clustering is its evaluation metrics. Jonyer et al. [213, 212] provides an objective evaluation metric that reflects the major requirements and tradeoffs of a good quality clustering. Yoneda et al.[282] present a method that can learn a graph structured representation from multivariate data where each node represents a cluster of data points and each edge represents the subset-superset relationship between clusters. The main idea of the method use formal concept analysis (FCA), which can extract hierarchical relationships between clusters based on the algebraic closedness property which leads to a directed graph representation.
5 Multidimensional Data Analysis
Multidimensional data analysis[249, 271, 248] allows the observation of data from various viewpoints.These data can most often be organized into meaningful hierarchies [272]. Multi-indicator models [250] are vey useful in the domain of multidimensional data analysis. Multiple indicators consist of two or more ”alternative” measures of the same concept. That is, a causal model contains alternative measures of the same thing. For example., two different ways of asking how satisfied you are with a service. Most of its familiar applications can be found in fields that involve the evaluation of subjective data. Many public opinion surveys employ this kind of data, usually measured among many thousands of individuals [49, 114]. Due to the complexity of multi-indicator systems[46], quantifying these indicators can be difficult and time expensive. Specifically, when dealing with nominal or ordinal multivariate data, a preliminary step is often scaling them with quantitative data. However, this can lead to inconsistent results[85] The most popular approach for quantifying multi-indicator data values is based on weighted-average or other aggregation techniques(additive[273], hierarchical[274] etc.) which is then used in a ranking procedure [39]. Ranking data represent a peculiar form of multivariate ordinal data taking values in the set of permutations [CITATION:TITLE:Mollica, PLMIX:AnRpackageformodelingandclusteringpartiallyrankeddata]. The mathematical aggregation of a set of individual indicators that measure multi-dimensional concepts leads to a composite indicator [275]. In the field of social sciences, composite indicators based on scoring or index functions are a commonly used method for ranking or measuring objects based on different properties[17]. A composite indicator may include several dimensions, where the dimensions represent different domains or aspects of the phenomenon being measured. For example, a composite indicator of well- being may cover dimensions such as income, employment, health and education. The problem of ranking subjects based on some ordinal scale is a fundamental issue in economics and social statistics. In addition, the validity and robustness of the results obtained from composite indices or other multidimensional data sources through aggregation has been challenged particularly due to the unavoidable aspect of subjectivity in their calculations. In addition, aggregative methods tend to oversimplify the complexity of a phenomenon and they also produce the same results for different situations [3, 5]. Moreover, such data often lack clear ordering criteria and many ambiguities arise when trying to compare individuals consistently[85]. In order to address these issues, many studies ([17, 31, 142, 87, 46, 49, 96, 10, 14, 48, 91, 61] etc. ) advocate the use of partially ordered set theory, which requires subjectivity only in the choice of the properties to be considered. The key motivation for advocating a poset oriented approach derives from the observation that classical aggregative methods of analysis may not clearly express the true complex nature of certain phenomena. Partial orders help to reveal why an object of interest holds a certain ranking position and how much it is subject to change if a composite indicator is upgraded [46]. This methodology is convenient in providing comparable indices or ranks of objects while reducing subjectivity to a minimum. The main advantages of managing multivariate data for ordering purposes based on the notion of posets for the purpose of anking are summarized as follows.
  • Partial order theory provides a technique to derive rankings in which ties can be considered, which avoids the need of a weighting of the indicators[62]. The horizontal arrangement of objects within a Hasse diagram and depicted by level structure gives a first approximation to a weak order (i.e. tied ranks are not excluded).
  • With partial orders there is no restriction to mapping indicators on a single scale.
  • Evaluation of phenomenon such as subjective well-being can be addressed in a consistent and effective way, using tools from partial order theory and overcoming the limitations of composite and counting paradigms[90].
  • Partial order theory and the related Hasse Diagram Technique provides tools for dealing with multidimensional systems of ordinal data which avoids the procedure of performing variable aggregation into composite indicators. As such, there is no necessity for converting ordinal scores into numerical values which can lead to inconsistency in the evaluation of the real phenomena under consideration [87, 10].
  • There is no requirement of “weighting” evaluation dimensions, to account for their different relevance when partial order theory is applied [87].
  • The evaluative process for ranking using partial order theory is advantageous in that the comparison of the objects of interest is done simultaneously for all indicators, without the need of any prior aggregation [63].
  • Partial order ranking offers a non-parametric method that neither includes any assumptions about linearity nor any assumptions about distribution properties [53].
  • Ranking by Hasse Diagram Technique has the advantage that it can be performed without any normative constraints. The theory of partially ordered sets and its graphical representation (Hasse diagrams) does not need any additional information to sort objects. Beyond sorting, many conclusions can be drawn from Hasse diagrams because they represent a well-defined mathematical structure [31, 27].
  • Posets are used to define the structure of comparabilities underlying multi-indicator systems [5]. Once the structure is defined, we can analyse it by using mathematical tools. The first goal of this analysis is the evaluation of scores to reduce the dimensionality of the phenomenon. We deal with quantitative variables so we analyse the resulting structure of comparabilities (the poset), without evaluating any difference or ratio to their values to get distances or proportions. Therefore, there is no loss of information.
We summarize how to obtain poset structures from subjective data set as follows. Let $k$ be the number of subjective indicators related to some observed phenomenon and let $n$ denote the number of individuals in the study such as a survey. Furthermore, let $\text{A}=\{q_{1},q_{2}...,q_{k}\}$ be such that each element of the set $A$ represents a different variable corresponding to a subjective indicator. Score the response information of each subjective indicator on an $m$ degree ordinal scale $(y_{1},y_{2}...,y_{m})$ which are ordered in such a way that $y_{1}<y_{2}<...<y_{m}.$ For each statistical unit associate the set of degrees on the variables. Each of the $n$ individual profiles is determined by the $m$ ordinal scores given in the set $B$ on each of the $k$ indicators in the set $A$. This is followed by pairwise comparison of all the $n$ individuals using their individual profiles corresponding to the ordinal scale in $B.$This evaluation procedure if not totally ordered leads to the poset structure which can be depicted by a Hasse graph of size $n.$ For instance, using this set up for a four degree scale, the profile $(y_{2},y_{3},y_{2},y_{3})$ is better than $(y_{4},y_{3},y_{3},y_{1}),$ but $(y_{3},y_{3},y_{4},y_{2})$ and $(y_{4},y_{3},y_{3},y_{2})$ are incomparable. Different types of methodology has been put forward which use poset theory for the analysis of multivariate datasets that have attributes which require multidimensional approach. Some of these are based on 1) fuzzy set approach [86, 85], partial order linear extension [87] 3) average rank [49] 4) poset theory with Adjusted Mazziotta–Pareto Index [4] 5) interpersonal comparability [155] 6) fuzzy-first order dominance [96] and so forth.
  • The fuzzy set approach was originally developed in the seminal work of Zadeh (1965)[251] as a class of sets that allows its members to have different degree of membership It can address uncertainty associated with vague concepts by allowing partial membership to a set. Fuzzy set approach has been used for the statistical evaluation of ordinal data associated to socio-economic phenomena in combination with posets in order to overcome the problems associated with the classical aggregative approach based on composite indicators[86, 85, 12]. The main feature of this approach is that the measurement level of the data is fully respected, avoiding any kind of not proper scaling.
  • In [87] a ranking method is developed based on the linear extensions from partial orders. Particulaly, it aims to solve the problem of ranking a finite collection of objects which can be represented as a cloud of points having a suite of indicator values corresponding to each member of the collection. It solves the problem of arbitrary assignment and aggregation of composite numerical score for each object , by viewing the relative positions as defining a partial ordering and any given pair of objects may not necessarily be comparable. The interval having the possible ranks assigned to each object can then be determined from the Hasse diagrams of the partial orders corresponding to the collection of all rankings that are compatible with the linear extensions from partial orders. In cases involving very large datasets which requires the enumeration a large number of linear extensions that can result to computational intractability, it is suggetsed in [142] to use the method of discrete Markov chain Monte Carlo for such intractable problems
  • Caperna & Boccuzzo[49] suggest a methodology for dealing with big complex data when some of the attributes require the measurement of complex concepts on ordinal and/or dichotomous scales. The method of average rank as a synthetic measure of a complex concept was motivated by the notion that poset theory is of paramount importance for dealing with ordinal data avoiding the use of scaling procedures. The main idea consist of a procedure for sampling units from a big population using a simple criterion to summarize the resulting values appropriately. This is followed by an application of the central limit theorem which makes it possible to compare the results obtained from different groups using statistical tests on the means. The Height of Groups by Sampling method is then used to compare the average rank among groups that are defined by one or more socio-demographic variables influencing the level of the measurable complex concept.
  • In [143] some procedures for n extracting insights on ordering properties that are embodied in multivariate datasets and applicable in configuring sets of indicators are reviewed.
  • The applied methodology [4] for synthesizing the evolution of multi-indicator systems over time is based on the partial order theory and the aggregative method of the Adjusted Mazziotta–Pareto Index, to one of the 15 sustainable development goals. In particular, the role of posets was to defne the structure of comparabilities underlying multi-indicators systems. This is then followed by a mathematical evaluation of scores to reduce the dimensionality of the phenomenon. Temporal posets are then obtained by merging the diferent posets. Furthermore, Aan embedded five level scale divided into( minimum, maximum, first, second and theird quartiles of the indicator values are also used to improve the quality of measurement.
  • Sen [155] propose a framework for interpersonal comparability in the context of aggregating the welfare measures of individuals. An aggregation relation between any pair of social states which is a quasi-ordering and identical with respect to the Pareto quasi-ordering under noncomparability, and is a complete ordering under unit comparability, and under full comparability are all considered.
  • Fattore & Maggino [90] outline a comprehensive procedure to address evaluation problems in a multidimensional ordinal setting. The procedure incorporates partial order theory for evaluation study that consist of identification of evaluation dimensions, addition of attribute relevance, threshold selection, computation of evaluation scores at the statistical unit and population level.
  • Fattore[92]: proposed a procedure for the assessment of multidimensional deprivation with ordinal data, based on partial order theory. The procedure focuses on achievement profiles and their comparabilities and incomparabilities The evaluation problem is treated in terms of multidimensional comparisons among profiles. The usefulness of the procedure is noted in its consideration for both vagueness and intensity of multidimensional deprivation and produces a complete set of synthetic indicators for the description of deprivation on statistical populations. In addition, t also allows for the introduction of information on attribute importance and it facilitates the handling missing data.
  • In [96] the procedure ”fuzzy-first order dominance ”(F-FOD), employs concepts and tools from partially ordered set theory and from fuzzy relational calculus with the goal of overcoming the main limitations of previously developed algorithms for FOD analysis. F-FOD is useful in that it produces full pairwise comparison matrices which allows for partial orderings and rankings of the statistical units to be derived from the input data.
The sorting approach using the Hasse Diagram Technique (HDT)[32], can be summarized as follows. Consider a dataset of objects to be sorted with respect to some criteria or property. Let m(i) be an appropriate attributes by which the objects are ordered. Furthermore, let $x$ and $y$ represent any two objects with $m(i,x)$ the value of the $i^{th}$ attribute of the object $x$($i=1,..,n$).Then $x\geq y$(which can mean that $x$ is evaluated better than or equal to $y$) if and only if $m(i,x)\geq m(i,y)$ for every $i=1,...,.,n$
6 Exploratory Data Analysis
In the field of data analysis, exploratory data analysis(EDA) is the preliminary investigation of data to discover patterns, spot anomalies, test hypotheses, check assumptions, relationships and insights. A major challenge in the exploratory data analysis phase is to identify noisy data which can significantly influence its outcome. Noisy data is often characterized by a high degree of variability which cannot be explained in the dataset. In partial orderings, there is the problem of data uncertainty as a result of data noise. Noise may cause changes in the overall orderings of objects. In order to address the issue of data noise, Bruggemann & Carlsen[44] construct a probability scheme which specifies a noise-model from which a distribution for noisy values explicit expressions are derived. This faciliattes a priori estimations which determines noise level effects with respect to the order relation between any pair of objects which are affected. As such, the expected level of noise perturbations can be quantified. Bruggemann & Carlsen [58] investigate the effects of data noise or uncertainties on the partial ordering of a series of objects by adopting a fuzzy approach such that partially ordered sets are obtained as a function of noise. Based on this, it becomes possible to identify the stability range in terms of noise where the original partial order is retained. Trend analysis can be a suitable method of exploratory data analysis that can be applied in the context of posets. As shown in [4], the evolution of Hasse diagrams can be analyzed over time for the synthesis of phenomenon from which a preliminary image of the temporal trend of the phenomenon can be derived. Missing data has been identified to be a problem in non linear trend analysis [270]. A poset methodology for handling missing values is enumerated in [92] using the case study of deprivation data. The pair-wise comparison of the objects of a poset can be a useful way of dealing witth missing data. The idea is that if there is no sufficient information to conmect a given object to another then there is still the possibility to have information to make comparison with other objects in the data set [17]. Formal concept analysis is particularly suited for exploratory data analysis[323, 317] because of its human-centeredness .
7 Descriptive Data Analysis
Descriptive analysis is one of the most crucial phases of statistical data analysis and data exploration that involves summarizing and describing the primary properties of a dataset. It also involves the process of using current and historical data to identify trends and relationships between variables. In addition, it is particularly useful for communicating the evolution of data and it uses trends as the basis for further analysis to drive decision-making. Poset theory has been used in a variety of contexts for data summarization and visualization. Hasse diagrams, serve the purpose of being the partial order graphical representation which conveys a considerable amount of information on the partial order structure of the data. For this, Hasse acyclic directed graphs of posets have been used for data visualization task that correspond to: causality[241], socioeconomic analysis [88], linear models and analysis of variance[242], experimental design [243], learning analytics[244], image analytics[245], medical data analysis [78], molecular structure prediction [183, 246], spatial analysis [121, 30] multidimensional data analytics [65, 46] etc. The principal aim of a visualization tool for multidimensional and partially ordered datasets is to provide a direct representation of the data structure, reducing its complexity but retaining the essential patterns in it [88]. Several formal data analytic methods have been developed that are based on the Hasse diagram. One of such methods is the Hasse Digram Technique(HDT) introduced by Brüggemann & Voigt[35] which is essentially an application of partial order theory based on a data matrix to analyze the structure of multivariate datasets, whenever a number of options can be characterized by multiple attributes (indicators). The Hasse diagram technique or use of posets can be useful in ranking problems that involve multi-criteria assessment[27] particularly when the objects from a dataset can be viewed relative to all others. There are several key motivations for the use of Hasse Diagram Technique for ranking problems as foolows:
  • it is parameter free and as such there are no assumptions needed about the statistical distributions of the attributes.
  • HDT makes the ranking process transparent.
  • The presentation by a Hasse diagram avoids the arbitrariness in constructing a ranking index [31].
  • HDT enables the extraction of structural information and the dimension of Hasse diagrams with respect to visualization of the ranking results [31].
  • HDT enables the visualization of parameters in sensitivity analysis [26].
  • Hasse diagrams allows a holistic view on all the objects which are to be ranked, without introducing artificial ranking indices [28].
In [32] an alternative approach using the Hasse Diagram Technique is proposed for sorting of chemicals with respect to their potential environmental hazard. In particular, in the Hasse Diagram Technique, the dominance of one chemical over another can be established whenever all attributes can simultaneously support the dominance. Additionally, [32] also proposes the average rank method which is based on the structure of the Hasse Duagram associated to the order relations encoded in the data matrix. A disadvantage of the HDT is the absence of aggregating functions, with weights as parameters, which would allow to remove conflictive descriptor values. In order to resolve that shortcoming, Brüggemann et al [158] developed METEOR(Method of evaluation by order theory) as an extended procedure which permits to solve the dilemma by obtaining a single high-ranked object, keeping the HDT transparency and allowing “researcher participation”. METEOR was subsequently applied for the evaluation of chemicals and environmental hazards in [159, 175] . Halfon et al.[107] suggest an innovative ranking procedure using a vectorial approach for partial ordering applicable to a variety of problems in environmental toxicology. The motivation to use the poset is that it can depict both not only situations of comparison amongs chemicals but also when chemicals are incomparable with respect to each other based on environmental hazard. The ranking is visualized with the help of computer programs displaying the Hasse diagrams that are valid for both large and small datasets comprising of 34 and 8 ranked chemicals respectively. In [191] the graphical technique called POSAC is developed for multivariate data analysis in a two-dimensional space such that observations and variables can be studied simultaneously. It works by mapping the rows (e.g., states) of a matrix in a way that maximizes the preservation of their partial order with similar states located in close proximity on the map. POSAC has been demonstrated to be useful in detecting the evolution of crime analytics in large geographical areas over time.
The set of chains and the antichains of a Hasse diagram represent important structural features that can be identified from their vertical and horizontal components respectively. Bruggemann[40] found that tripartite graphs are useful in the interpretation of complex data sets by clarifying the role of indicators causing incomparabilities which are associated with the horizontal components of Hasse diagrams. In the study of the evaluation of analytic performance using partial ordering in [57], it is demonstrated that the presence of incomparabilities (being an inherent feature in partial order ranking) may disclose important characteristics. It is also shown that summary statistics of the absolute z-score, absolute skewnesss and the standard deviation from calcualtions using data on analytical performnce can be visualised on a Hasse diagram to provide a holistic view of the results.
The well-established use of lattice theory is in formal concept analysis (FCA) which provides a conceptual framework for structuring, analyzing and visualizing data, in order to make them more understandable [324]. In particular, the concept lattice with its diagram graph allows the visualization and summarization of data in a more concise representation. The main point of FCA which is of interest here is the “symmetric” view on objects and properties, whereas a Hasse diagram, as shown in only shows the order relations focusing on the objects [40].
A limitation of a Hasse diagram as a representation of a poset is as a result of its complexity when there exist a large number of edges and vertices. In this case, they are not easy to interpret as the number of vertices increases and, as noticed and they do not also provide any metric information when this is available in the original data. Cluster analysis, on the other hand, reduces the complexity of the data, but it is not designed to preserve information on comparabilities and incomparabilities. Bruggemann & Carlsen [247] combine Hasse diagrams and cluster analysis in a complexity reduction process, producing a visual output that allows the end users to jointly grasp the partial order and the metric structure of the data. This visualization process involved the following steps [88]:
  1. 1.
    Reducing dataset complexity through a clustering process based on a SOM(Self-organizing map).
  2. 2.
    Building a classical Hasse diagram on the population of clusters, that is on SOM codebooks.
  3. 3.
    Visually adding information on statistical units and clusters (particularly, information pertaining to the value of the covariates)
The use of cluster analysis and principal component analysis with HDT may be helpful in obtaining a statistically relevant data representation and in avoiding insignificant numerical differences of the attributes, which in turn would lead to insignificant comparabilities and incomparabilities and thus to very complex Hasse diagrams.[31]]
Large data sets may be difficult to deal with computationally in FCA. More especially, it is the number of formal concepts derived from a data set that is the key factor in determining if a concept lattice will be useful as a visualisation. Andrews& Orphanides [6] suggests that understandable results can be obtainable from existing data sources where this would not normally have been the case whereby formal contexts that would normally be intractable for visualisation have been processed by firstly, focusing on information of interest and secondly, reducing ‘noise’ in the context, thus revealing readable lattices that lucidly portray conceptual meaning in large data sets.
8 Applications
In this section, we summarize the data sets, evaluation methods, software and algorithmic solutions for the analysis of partially orderd data. We detail practical applications of posets in various domains of machine learning and data analysis. Since as partially ordered data and structures are ubiquitous, posets have a wide variety of applications. Although the applications are very broad, a significant number of its applications pertains to data-driven decisions in problems of sustainable development[5, 109, 4]. In this case, the main examples pertain to the analysis of complex and multidimensional systems of ordinal data and to problems of multi-criteria decision making which are relevant in social and environmental sciences. A comprehensive review of the theory behind the applications of posets in socio-econics and environmental sciences is found in [94]. The application of partial order methodology can sometimes be used as an interim process before other tools will be applied is emphasized[63]. Posets and other methods arising from lattice theory are generally useful in comparative evaluation processes [27]. It is more appropriate to use poset methodology for performing comparative evaluation than using statistical tools since as decisions on binary choices based on comparisons by applying some criteria are naturally based on order theory. Furthermore, ordering based on posets can be visualised and additional results can be derived from the mathematical theory of relations between objects. Order relations also facilitate the identification and evaluation of most relevant objects in a study as well as deriving the sequence of the importance of a criteria to perform a ranking process for ordinal data. The Hasse diagram technique remains a powerful tool to perform comparative evaluations [30]. There are several techniques for quickly identifying the minimal and maximal elements of a poset even in complex scenarios which are of fundamental interest whenever priority setting procedures are to be performed. Partially ordered sets often comprise a big amount of information on the degree of dominance among their elements and proper tools can be employed to extract and turn it into synthetic scores and rankings.
8.1 Environmental Chemistry and Toxicology
The concept of partially ordered sets and their visualisation by Hasse diagrams turns out to be very useful in many applications of environmental concern where evaluative considerations are important [107]. Environmental data have been extensively studied by Hasse Diagram Technique in [29, 30, 31, 33] with the goal to identify major relationships both among sediment samples (objects) and degradation indices (attributes). Further investigations using formal concept analysis[7] with the aim to highlight interaction among hygienic compounds and a synergism between the two toxicity tests applied to the sediment samples Lake Ontario. Hasse diagram technique[30] supports the visualization of more than two-dimensional problems to identify pollution patterns by characterizing geographical regions by their pollution with chemicals.and to suggest priority regions for further examination. The pollution pattern responsible for the location of the regions within a Hasse diagram, helps to look for remediation strategies. In alternative approach, the pollution pattern of 59 regions of Baden Wuerttemberg Germany using a methodology that combined hierarchical clustering analysis(HCA) and Hasse Diagram Technique [240] . The HCA was used for classification of the objects in the set in order to find similarity classes resulting to reduced set of object repreentatives thereby making the Hasse diagrams for analysis depicting the network structure more easily understandable. Posets can be very useful when ranking environmental chemicals by multicriteria analysis [36, 277, 9]. Posets in this case provides a solid formal framework for the ranking of objects without assigning a common scale and/or weights to the criteria. In the prediction of toxicity levels of chlorobenzene in environmental chemical polychlorinated hydrocarbons, a scheme[252] is developed based on poset theory which embeds an overall reaction network
8.2 Socio-economic Analysis
Partially ordered data are prevalent in many branches of the social and behavioral sciences. This has been driven by the fact that partial order theory employed in the applied sciences overcome the intrinsic disadvantage hidden in aggregation, if a multiple attribute system is available[10]. A simple example is response categories Agree, Neutral, Disagree, and Don’t Know, of which the first three can be ordered and the last forms a category of its own. This type of data can be derived from a wide range of applications covering different topics: multidimensional poverty, economic development, inequality measurement and many more. Partially ordered set theory has been shown to be particularly suitable to address multi-criteria decision problems since it shows where multidimensional indicator data values are expressing a conflict which can then be easily identified and resolved by appropriate corrective action [65]. The use of posets with quantitative data reduces the set of operations and choices to be made in order to synthesize indicators (normalization, aggregation), even if they are the natural representation of multidimensional ordinal data [92, 93]. Using this approach, it is possible not only to investigate the nature of the phenomenon in more detail, but also to help policy-makers in their assessments. Infact, it can be considered an efective policy tool, aimed at promoting the identifcation of disadvantaged and under-developed profiles and contexts [5]. The application of the poset-based method provides an understanding of the complexity of the evolution of a social phenomena in terms of temporal trends and comparisons between regional structures in a countrywide basis using a multideimensional data source for the process [4, 49].Incomparability captures the existence of intrinsically different “shapes” of it, so providing a more realistic picture of the phenomenon under investigation. In this respect, the impossibility to compare the profiles of different cities, due to the existence of dimensions where they perform in conflicting ways, reveals the irreducible complexity of sustainability. In census data on deprivation within regions of a country[12, 85, 86] posets has been applied since they can account for incomparabilities, which are at the basis of deprivation complexity. Partial order theory has been extensively applied to the construction of synthetic indicators in various socio-economic contexts [4, 5, 8, 3, 18, 77, 39, 59, 60, 74, 91, 93, 95, 97, 254, 255]. In the area of crime analytics[48] POSAC has been used to study structure of criminal networks in terms of the longevity, the biographical data the network sizze and so forththe The main finding is that POSAC examination of measures derived from SNA reveals[There are other citations on CRIME and Posets INclude]
8.3 Neurocognition Modelling
Finite partially ordered sets are natural models for cognition, as it is reasonable to assume that some cognitive states have higher levels of functionality than others[165, 166, 164, 169, 170, 170]. Finite partially ordered classification models are useful for many statistical applications, including cognitive modelling. A data analytic framework for implementing latent finite partially ordered classification models is introduced which proposes such tools which is useful cognitive applications that are latent and complex. Poset models are flexible and can become quite rich and complex, enabling them to be effective models for describing response phenomena from test educational test data or neuropsychological assessment data[165, 167, 173]. An objective of neuropsychological assessment is to determine differences in cognitive functioning in clinical settings. Carr et al. [50] used poset classification models of neuropsychological test data to classify samples to detailed cognitive profiles using ADNI2 and AIBL data. In the risk of disease progression of Alzheimer’s disease for individiuals with mild cognitive impairment (MCI), Tatsuoka et al. [171] suggest that poset-based modeling methods may be useful for providing more precise classification of cognitive subgroups among MCI for imaging and genetics studies, and for developing more efficient and focused cognitive test batteries. An order structure arises naturally with skill profiles. The flexibility to not necessarily assume that one state is greater than another is an appealing feature of posets[169]. In addition, posets are comprised of states, into which cases are classified, that are associated with distinct patterns of attribute strengths and weaknesses [120]. Posets have several advantages over conventional statistical methods for handling large numbers of polyfactorial neuropsychological test variables such as the ability to mimic the expert judgment of a clinical neuropsychologist for each case in a large sample. They are efficient since valid conclusions can be drawn based on relatively few measures and classification of large samples can be accomplished rapidly [119]. However, Jaeger et al. [119] noted the limitation in the poset model applied to large conventional neuropsychological test dataset is that it is restricted to a selected set of attributes, while additional important distinctions remain to be tested.
8.4 Other
The application of posets and lattice models is not limited to the afore- mentioned domains and tasks for data analysis. There have been explorations of applying posets to a variety of problems such as hypothesis management in large scale research projects [104]; adaptive testing for cognitive assessment [172]; machine Learning of reviewer’s preferences on papers [268], similarity learning use pose estimations [20], and many more data driven applications.
8.5 Some Selected Datasets
We classify the data sets into: Sustainable development dataset; Computer Vision Dataset; Natural Language
Computer Vision Datasets
  • In the study of pose estimation in [20] for the study of visual similarity learning using posets, three data sets were considered: The Olympic Sports dataset [331], The Leeds Sports Pose(LSP)[330], MPII Pose [329].
  • Micro-video dataset [311]. It is used for the learning problem on hypergraph partial order [99].
Natural Language Datasets:
  • The Compositional Freebase Questions (CFQ)(Keysers et al., 2020)[332] is a dataset that is specifically designed to measure compositional generalization. It is used for the study of hierarchical poset decoding for compositional generalization in language [105].
  • Universal Dependencies (UD) corpora [333]. It is used in the learning algorithm for generating a surface word order for a sentence given its dependency tree [83].
Sustainanble Development Data:
Sustainability data have been subdivided in two datasets, the first comprising socio-economic indicators and the second comprising environmental data.
Socio-economic data
  • Equitable and sustainable well-being data [334]. Used in [4, 3] for multidimensional data analysis in the context of sustainable development.
  • Service Performance Dataset [335]. Used for multidimensional data analysis in [10].
Environmental data
  • A battery of biochemical, microbiological and bioassay tests were used to identify degraded or degrading sediments and waters [29, 31, 7]. The dat set is available in [82].
  • Collection of toxicity data [252].
8.6 Selected software algorithms and packages
Machine learning and Deep learning
  • Python implementation of deep visual similarity learning using posets [336]. Based on [20].
  • [83] Given an edge-weighted poset, the algorithm in [336] constructs a total order such that nodes with smallest weights are adjacent. The algorithm works by attempting to order a set of words as closely as possible to their original surface realization in the UD corpus. Due to the fact that words may repeat in the sentence, each order is instead represented by a list of integers, and it is these lists of integers which are compared [336]. For example, assuming a target reference order of [1,2,3] for the red horse, the generated order of red the horse would be [2,1,3].
  • Causal Structure Discovery Algorithm-Greedy Sparsest Poset (GSPo) , python implementation in [337]. Used for causal structure learning in the presence of latent variables [22].
Multidimensional data analyis:
  • The PARSEC is an An R package for poset-based evaluation of multidimen-sional poverty which was introduced in [11]. Its main goal is to provide socio-economic scholars with an integrated set of elementary functions for multidimensional poverty evaluation, based on ordinal information. The package is organized in four main parts
    1. 1.
      Data management.
    2. 2.
      Basic poset analysis.
    3. 3.
      Poset-based evaluation.
    4. 4.
      OPHI counting approach.
  • Poset algorithm for multidimensional data analysis [13].
  • Self Organizing Map(SOM)Self-organizing map algorithm[286], which is one of the most used tools for non-linear dimensionality reduction and pattern recognitio, in multidimensional data analysis. It has been used in:[12].
  • In [13] illustrate a non-aggregative algorithm new way of constructing sustainability synthetic indicators on multi-indicators systems.
  • A novel application of partial order theory for object rank correlation analysis of multiple variables is presented by the new software package PO Correlation [162] . The design is made transparent for rank correlation analysis by a detailed mapping of the rank relations between all objects.
  • The software WHASSE is applied for chemical monitoring data analysis [32, 34], medical data analysis [78].
  • PLMIX[139] an R package for modeling and clustering partially ranked data.
  • PyHasse Software [42, 51, 127] is for the pupose of the ordinal analysis of data matrices, identification and analysis of partial order relations as well as computing ranks.. It is used in [179] for the analysis of environmental data .
9 Future Research Directions
Posets have proven their usefulness in a broad range of applications. But there still remains many more data centric applications where existing and new methods of learning and analyzing partially ordered data can be relevant in generating novel results. In this section we suggest four areas.
Learning Algorithms A major challenge in data science is the identification of geometric structure in high-dimensional data. The structural understanding of data is very relevant for designing efficient algorithms for optimization and machine learning. Classically, the structure of data has been studied under Euclidean assumptions. The fundamental representation of any feature for any machine learning model is the vector and its multidimensional generalization which is the tensor. As such many machine learning algorithms and data analyis pipelines have been developed with the aim of computing vectors or matrices of real numbers. There are currently a wide range of well-studied tools and algorithms that assume such structure. However, many scientific fields study data with an underlying structure that can only be represented in non-Euclidean space. Such data can be found in graphs and many kinds of manifolds. There is currently growing research on developing machine learning models that work with data representations for features and embeddings, and or model parameters in the non Euclidean domains. Posets as a class of directed acyclic graphs have also begun to recieved attention with regards to developing appropriate learning algorithms. Recently, Wendler [180] introduced new methods for the fourier-sparse learning on data indexed by lattices and posets. The graph neural network class of algorithms also works on posets as specified in section 3. But there is not yet widespread use of graph neural networks relative to the number of its applications based on posets and lattices. The success of machine learning algorithms generally depends on data representation since as different representations encode different explanatory factors of variation behind the data [283]. There is enormous scope to develop new robust and efficient algorithms for learning the key structural characteristics of data represented as lattices and posets.
Posets and Topological Data Analysis Topological data analysis (TDA)[325] is a recent and fast-growing field in applied mathematics which provides a set of new topological and geometric tools to infer relevant features for complex data. There is a growing body of literature which connects posets to topology data analysis. This current article did not take into consideration the posets that appear in topological data analysis. There are many new applications which can arise from the study of toplogical analysis of partially ordered data.
Machine Learning and Formal Concept Analysis Several key aspects of machine learning in the context of formal concept analysis was introduced in section 3.9. However, the scope of applications is not exhaustive especially for ontology learning using formal concept analysis techniques. A more indepth review of existing and new models for machine learning using formal concept approaches can be useful in highlighting latest results.
Learning to rank with partially ordered data Learning to rank refers to machine learning techniques for training the model in a ranking task. The prediction of structured outputs in general and rankings in particular has attracted considerable attention in machine learning in recent years, and different types of ranking problems have already been studied. There has been many ranking problems introduced in the context of data analysis of partially ordered data in sections 5 and 7. There is a great potential of introducing existing and new techniques of learning to rank for ranking problems with datasets from which can be derived a partial order structure such as big data for sustainable development etc. In particular, it can be useful to implement solutions from the various graph based[339, 338] approches of learning to rank for this purpose or developing new learning to rank algorithms for partially ordered data.
10 Conclusion
This article can be considered the first exhaustive review summarizing the applications of posets in relation to data analysis and machine learning. The vast use of poset theory in many domains of machine learning and data analysis demonstrates their relevance for a long time to come. We have examined the growing adoption of poset methodology in several areas of machine learning and deep learning. Posets theory was also systematically categorized into multidimensional, descriptive and explorative data analysis from many perspectives. The practical applications of posets was drawn from a broad range of domains. The main contributions of this work is to highlight the deep connections poset theory has with data analysis and machine learning. It is expected that this article will provide a basic and in-depth understanding of the applications of posets in data science and the discovery of new directions for future research on the subject.